OpenStudy (anonymous):

Find the quotient. x+4 / x^2 divided by 2/x a. 2x / x+4 b. 2x^2 / x+4 c. x+4 / 2x d. x+4 / 2x^2 help me please. :) <3

OpenStudy (ipwnbunnies):

\[\frac{x+4}{x^2} \div \frac{2}{x}\] When dividing rational expression. Keep the first. Change the sign to multiplication. Flip the second fraction. \[\frac{x+4}{x^2} * \frac{x}{2}\]

OpenStudy (ipwnbunnies):

It's not D. ;-;

OpenStudy (aric200):

I was eliminating D

OpenStudy (aric200):

Ty for letting me finish

OpenStudy (ipwnbunnies):

LOL

OpenStudy (anonymous):

\[\frac{ x+4 }{ x^2 }\div \frac{ 2 }{ x }=\frac{ x+4 }{ x^2 }\times \frac{ x }{ 2 }=?\]

OpenStudy (aric200):

Just gonna casually go now

OpenStudy (aric200):

I simplified okay... Leave me alone

OpenStudy (anonymous):

B? :)

OpenStudy (aric200):

F!!!!!!!

OpenStudy (ipwnbunnies):

.-. You can simplify it. There's an x in the numerator, an x^2 in the denominator. The 'x' in the numerator cancels with one in the denominator, leaving you with: \[\frac{x+4}{2x}\]

OpenStudy (anonymous):

c okay thanks :D

OpenStudy (aric200):

Welcome

OpenStudy (aric200):

I eliminated D okay So...